# DiffSensei: Bridging Multi-Modal LLMs and Diffusion Models for Customized Manga Generation

### v1

@@ Abstract
Story visualization, the task of creating visual narratives from textual descriptions, has seen progress with text-to-image generation models. However, these models often lack effective control over character appearances and interactions, particularly in multi-character scenes. To address these limitations, we propose a new task: customized manga generation and introduce DiffSensei, an innovative framework specifically designed for generating manga with dynamic multi-character control. DiffSensei integrates a diffusion-based image generator with a multimodal large language model (MLLM) that acts as a text-compatible identity adapter. Our approach employs masked cross-attention to seamlessly incorporate character features, enabling precise layout control without direct pixel transfer. Additionally, the MLLM-based adapter adjusts character features to align with panel-specific text cues, allowing flexible adjustments in character expressions, poses, and actions. We also introduce MangaZero, a large-scale dataset tailored to this task, containing 43,264 manga pages and 427,147 annotated panels, supporting the visualization of varied character interactions and movements across sequential frames. Extensive experiments demonstrate that DiffSensei outperforms existing models, marking a significant advancement in manga generation by enabling text-adaptable character customization. The code, model, and dataset will be open-sourced to the community.  [^1] The work is done in Shanghai AI Laboratory.
Figure 1: Results of DiffSensei. (a) Customized manga generation with controllable character images, panel captions, and layout conditions. Our DiffSensei successfully generates detailed character expressions and states following the panel captions. (b) Manga creation for real human images. The dialogues are post-edited by humans. The continuation is in the Appendix Fig. 7. We strongly recommend that the readers see the Appendix for more comprehensive results. Manga reading order: Right to left. Top to bottom.
@@ 2 Related Work
Story visualization. Story visualization, the process of generating visual narratives based on given stories, is rapidly evolving. Many approaches can generate consistent image sequences derived from story content [52, 36, 25, 21, 5, 45]. Despite recent advancements, the field faces significant limitations. Most existing methods generate story images solely from text and image-level prompts [52, 25, 21, 45], which restricts control over individual characters. This limited control over characters reduces the flexibility and depth of story visualization. A key factor is that current training datasets [15, 18, 10, 21, 45] lack character-specific annotations. In response to the data limit, recent works [36, 5] explore multi-character control using training-free methods that leverage existing subject preservation techniques, such as IP-Adapter [46]. Other works [7, 29, 9, 46, 4, 38, 35, 19, 43, 12, 11] try to train diffusion models for a multi-character customized generation. However, these approaches often result in a "copy-pasting" effect, restricting the diversity of expressions and actions needed for dynamic storytelling. For training-free methods, combining multiple models can significantly slow down inference speeds. To address these challenges, we first introduce a large-scale manga generation dataset with finely curated character annotations and then develop a novel framework utilizing diffusion models and MLLMs that enables the dynamic generation of manga panels.
@@
In summary, we propose the first large-scale dataset for customized manga generation, along with a novel framework tailored for this task. It utilizes an MLLM as a character feature adapter, significantly improving the text compatibility of character personalization.
@@ 5 Experiments
In this section, we conduct a thorough evaluation of DiffSensei and compare it with baseline models.
@@ Appendix D Ablation Study
Rate of character image sourced from target panel. Thanks to the character annotations, we can identify the same character across multiple panels, which allows us to capture different appearances, poses, and motions. During training, we sample a source character using its character ID, which remains consistent throughout the page. However, we find that using a completely random sample made training difficult to converge, likely due to the artistic exaggerations often found in manga. These exaggerations make it challenging for the model to learn consistent character representations when the character’s appearance changes significantly across panels. To address this, we introduced an alternative sampling strategy where, at a set rate, the source character is sampled directly from the target image itself. This helps the model learn identity preservation more effectively. An ablation study examining the impact of the target character sampling rate is presented in Tab. 4(a). The results indicate that a higher sampling rate leads to a decrease in text compatibility, as the model becomes overly focused on replicating the input character image. On the other hand, a lower sampling rate makes it difficult to converge the training, which also negatively impacts the metrics. Ultimately, we choose a sampling rate of 0.5 during training to provide a balanced performance.

### v2

@@ Abstract
Story visualization, the task of creating visual narratives from textual descriptions, has seen progress with text-to-image generation models. However, these models often lack effective control over character appearances and interactions, particularly in multi-character scenes. To address these limitations, we propose a new task: customized manga generation and introduce DiffSensei, an innovative framework specifically designed for generating manga with dynamic multi-character control. DiffSensei integrates a diffusion-based image generator with a multimodal large language model (MLLM) that acts as a text-compatible identity adapter. Our approach employs masked cross-attention to seamlessly incorporate character features, enabling precise layout control without direct pixel transfer. Additionally, the MLLM-based adapter adjusts character features to align with panel-specific text cues, allowing flexible adjustments in character expressions, poses, and actions. We also introduce MangaZero, a large-scale dataset tailored to this task, containing 43,264 manga pages and 427,147 annotated panels, supporting the visualization of varied character interactions and movements across sequential frames. Extensive experiments demonstrate that DiffSensei outperforms existing models, marking a significant advancement in manga generation by enabling text-adaptable character customization. The code, model, and dataset are open-sourced to the community.  [^1] The work is done in Shanghai AI Laboratory. Corresponding to: Xiangtai Li, Jingbo Wang.
Figure 1: Results of DiffSensei. (a) Customized manga generation with controllable character images, panel captions, and layout conditions. Our DiffSensei successfully generates detailed character expressions and states following the panel captions. (b) Manga creation for real human images. The dialogues are post-edited by humans. The continuation is in the Appendix. We strongly recommend that the readers see the Appendix for more comprehensive results. Manga reading order: Right to left. Top to bottom.
@@ 2 Related Work
Story visualization. Story visualization, the process of generating visual narratives based on given stories, is rapidly evolving. Many approaches can generate consistent image sequences derived from story content [53, 37, 25, 21, 5, 46]. Despite recent advancements, the field faces significant limitations. Most existing methods generate story images solely from text and image-level prompts [53, 25, 21, 46], which restricts control over individual characters. This limited control over characters reduces the flexibility and depth of story visualization. A key factor is that current training datasets [15, 18, 10, 21, 46] lack character-specific annotations. In response to the data limit, recent works [37, 5] explore multi-character control using training-free methods that leverage existing subject preservation techniques, such as IP-Adapter [47]. Other works [7, 29, 9, 47, 4, 39, 36, 19, 44, 12, 11, 32] try to train diffusion models for a multi-character customized generation. However, these approaches often result in a "copy-pasting" effect, restricting the diversity of expressions and actions needed for dynamic storytelling. For training-free methods, combining multiple models can significantly slow down inference speeds. To address these challenges, we first introduce a large-scale manga generation dataset with finely curated character annotations and then develop a novel framework utilizing diffusion models and MLLMs that enables the dynamic generation of manga panels.
@@ 5 Experiments
In this section, we thoroughly evaluate DiffSensei and compare it with baseline models.
@@ 6 Conclusion
Acknowledgement. This work is supported by the National Key Research and Development Program of China (No. 2023YFC3807600).
@@ Appendix B More Qualitative Comparison Results
Figure 9: Comparison with ground truth. The results generated by DiffSensei show remarkable consistency and aesthetics compared to ground truth manga panels.
@@
Moreover, we show the comparison with ground truth in Fig. 9. The result highlights the remarkable visual quality and character consistency of DiffSensei. DiffSensei achieves close or even better character image alignment than ground truth.
@@ Appendix C Implementation Details
Human evaluation details.
Figure 10: Human evaluation interface.
For human evaluation, we invited 15 general users from diverse backgrounds, ages, and genders to assess 10 anonymized manga page sets each. Evaluators reviewed the pages from different aspects, and their ratings were averaged to compute the preference probability across methods. The interface is shown in Fig. 10.
@@ Appendix D Ablation Study
Rate of character image sourced from target panel. Thanks to the character annotations, we can identify the same character across multiple panels, which allows us to capture different appearances, poses, and motions. During training, we sample a source character using its character ID, which remains consistent throughout the page. However, we find that using a completely random sample made training difficult to converge, likely due to the artistic exaggerations often found in manga. These exaggerations make it challenging for the model to learn consistent character representations when the character’s appearance changes significantly across panels. To address this, we introduced an alternative sampling strategy where, at a set rate, the source character is sampled directly from the target image itself. This helps the model learn identity preservation more effectively. An ablation study examining the impact of the target character sampling rate is presented in Tab. 4(a). The results indicate that a higher sampling rate decreases text compatibility, as the model becomes overly focused on replicating the input character image. On the other hand, a lower sampling rate makes it difficult to converge the training, which also negatively impacts the metrics. Ultimately, we choose a sampling rate of 0.5 during training to provide a balanced performance.
Human evaluation of the MLLM module. We conducted a human evaluation on the MLLM module. The results are presented in Fig. 12. The full model with MLLM demonstrates a significant advantage, particularly in terms of text alignment and overall preference. The results indicate that the MLLM module enhances both the visual appeal and text compatibility of manga generation.